OpenStudy (anonymous):

Do both a^n + a^n and a^n + a^2 equal a^2n?

OpenStudy (anonymous):

a^n +a^n =2a^n and a^n + a^2=?

OpenStudy (anonymous):

No, the first.\[a^n + a^n = 2a^n\]The second is harder to simplify but can be changed to\[a^n + a^2 = a^2(a^n-^2 +1)\]

OpenStudy (anonymous):

both are not equal to a^2

OpenStudy (anonymous):

if n=2 then it is possible